calculate the bond equivilent yield of this bond (quoted rate): Answer should be in 3 decimal places and whole numbers: ex: 6.125 not 6.125% or 0.06125 Price $89.5 Coupon: 5.5% Maturity Date: 9/15/2026 Today's Date: 9/15/2018 Compounding/Payments: Semi-Annual
Answer #1

Using financial calculator

N=8*2=16

PV=-89.5

FV=100

PMT=5.5%*100/2=2.75

CPT I/Y=3.626%

Hence, bond equivalent yield=3.626%*2=7.253%